What is the maximum distance you could place an Omega Pressure transmitter Model PX119-030G from the current measuring instrument if you are using 24 AWG wire (25Ω/1000ft) and using a 18 VDC power supply?
Answer #1

Considering 5% maximum voltage drop as standard

5% of 18V = 0.9 V

Also max current used by omega transmitter = 20mA

(from data sheet of Model PX119-030G)

So , Voltage drop =0.9 = I * R = 20mA *R

So , R= 0.9/ 20mA = 45 ohms

So 25 ohms have 1000ft length

So 45 ohm will have length = 45*1000/25= 1800 feet

So maximum distance we can use is 1800 feet!!